# Bayesian Analysis of High Dimensional Vector Error Correction Model

### 28 Dec 2023 v1

@@ 1 Introduction
Let $Y$ denote a $p$-dimensional multivariate time series, with $Y_{t}\in\mathbb{R}^{p}$ being observed at time $t\in[T]$, and $Y_{t,j}\in\mathbb{R}$ denoting observation of $j$-th variable at time $t$. We write $Y(\cdot,j)$ to denote the $j$-th variable of vector $Y$, which is a univariate time series. Stationarity of $Y(\cdot,j)$ is considered as the second-order stationarity. Precisely, we say $Y(\cdot,j)$ is stationary if the first moment being finite and independent of time and that the auto-covariance function being finite and independent of time. We denote $Y(\cdot,j)\sim I(0)$ in this scenario. When $Y(\cdot,j)\sim I(0)$ for all $j\in[p]$, we write $Y\sim I(0)$. When $Y(\cdot,j)$ is non-stationary but when $\Delta Y(\cdot,j)$ defined by $\Delta Y(t,j):=Y(t,j)-Y(t-1,j)$ is stationary, we write $Y(\cdot,j)\sim I(1)$. Particularly, when $Y(\cdot,j)\sim I(1)$ for all $j\in[p]$, we write $Y\sim I(1)$.
@@ 1.1 VECM and a motivation for High Dimensional VECM
The idea of high dimensional VECM falls in the trendy literature surrounding ‘big data’, in which we have a large amount of variables of interest ($p$) and a substantial amount of observations taken over time ($T$). In a traditional statistical setting, we require more observations than unknown parameters — in the setting of Equation 2 we need $T>p^{2}+\frac{p(p+1)}{2}$. However, that often means we need a greater amount of observations than the data actually have, in order to estimate the equation. This is particularly concerning in light of change-points or other potential changes of the data over time, which naturally constrains the $T$ available. High dimensional VECM is therefore motivated in the regime where $p,T$ are large but $T\leq p^{2}$. Certainly, for estimating the non-sparse coordinates of the matrix $\Pi$, it is reasonable to impose a feasibility assumption where $T>pr$.
@@ A QR decomposition is then proposed, taking the form of
$\tilde{\Pi}=(AB^{\intercal})(BB^{\intercal})^{-1}$ (4)
where $\tilde{R}\in UT(p)$ and $\tilde{S}\in\mathbb{R}^{p\times p}$ is orthogonal. Consequentially $\tilde{\Pi}=\tilde{R}^{\intercal}\tilde{S}^{\intercal}$ where $\tilde{R}^{\intercal}$ is a lower triangular matrix. [5] proceed with group-lasso regression, however, we opted for a different route to investigate the true rank of $\Pi$.
The SSL literature mainly concerns scalar response multivariate-regression, i.e. the type taking the form of $y_{i}=x_{i}^{\intercal}\beta+\varepsilon_{i}$ for each index $i$, with $x_{i}$ and $\beta$ being $p-$dimensional vectors and $y_{i}$ being an one-dimensional observation, while $\varepsilon_{i}$ is an iid draw from a Gaussian distribution with zero mean and unknown variance. $\beta_{0}$ is denoted as the true value of the $\beta$ parameter. SSL can be thought of as a mixing distribution being imposed on the parameter $\beta$, conditional on certain hierarchical Bayes parameters $\gamma$, which is then conditionally Bernoulli based on $\theta$, which denotes prior expected fraction of non-zero entries of the true $\beta_{0}$. A posterior mode estimator then aims to maximise the posterior distribution $\pi(\beta|\theta)$. Write $\hat{\beta}=\operatorname*{argmax}\pi(\beta|\theta)$. In [7], a Beta distribution on $\theta$ was proposed to enable further inference on the distribution of the estimator conditional on $\theta$, denoted $\pi(\hat{\beta}|\theta)$, and thereafter the conditional expectation $\mathbb{E}[\theta|\hat{\beta}]$. In [8], further analyses were drawn on the asymptotic behaviour of $\hat{\beta}$ such as bounding the limiting probability of $\sup_{\beta_{0}}\mathbb{E}_{\beta_{0}}\left[\mathbb{P}[\hat{\beta}:||\hat{\beta}-\beta_{0}||_{1}>\varepsilon]\right]$ with $\varepsilon$ depending on the number of observation $n$ and features dimension $p$.
@@ 2 Methodology
2.1 Hierarchical Bayes Modelling of SSL on VECM
@@
Let $X\in\mathbb{R}$ be a random variable. A Spike-and-Slab Lasso (SSL) pior on $X$ is defined as
@@
We write $X|\gamma\sim SSL(\gamma,\lambda_{0},\lambda_{1})$
We may interpret $X|\gamma$ as a mixture of two Laplace distribution, with $\psi_{0}$ being spiky at around 0 and $\psi_{1}$ less spiky.
Recall the Equation 3 and Equation 4 composition of the pre-estimate of $\Pi$. Here, we attribute the following definition of SSL distribution on the decomposed VECM. We use the notation $A_{t}:=\Delta Y_{t}$ and $B_{t}:=\tilde{S}^{T}Y_{t-1}$ and centralise them into $\tilde{A}_{t},\tilde{B}_{t}$ respectively (see subsection A.1 for granular details of data processing).
Definition 2 (SSL distribution on decomposed VECM)
@@
Let $\tilde{A}_{t},\tilde{B}_{t}\in\mathbb{R}^{p}$ for all $t\in[T]$, and define $e_{t}:UT(p)\to\mathbb{R}^{p}$ by $e_{t}(R)=\tilde{A}_{t}-R^{\intercal}\tilde{B}_{t}$ for all $t$. An SSL distribution on decomposed VECM takes the form of
$\tilde{A}_{t}|\tilde{B}_{t},R,\sigma^{2}$ $\sim N(R^{\intercal}B_{t},\sigma^{2}I_{p}),\ \ \pi(\sigma^{2})\propto\sigma^{-2}$ $\forall t\in[T]$ (7)
$R(\cdot,j)|\gamma_{j}$ $\sim SSL(\gamma_{j},\lambda_{0,j},\lambda_{1})$ $\forall j\in[p]$ (8)
$\gamma_{j}|\theta_{j}$ $\sim Bernoulli(\theta_{j})$ $\forall j\in[p]$ (9)
$\theta_{j}$ $\sim{\cal B}(a_{j},b_{j})$ $\forall j\in[p]$ (10)
Under this setting, we have the Bayes Estimator
$(\hat{R}^{BE},\hat{\sigma}^{BE})=$ $\operatorname*{argmin}_{R\in UT(p),\sigma\in\mathbb{R}_{>0}}\sum_{t\in[T]}\frac{||{e_{t}}(R)||_{2}^{2}}{2\sigma^{2}}+(n+2)\log(\sigma)+\text{pen}(R,\hat{\theta}^{BE})$ (11)
@@
The first two terms of Equation 11 are due to Gaussian assumption (Equation 7) with Jeffery prior on the variance. The penalty term $\text{pen}(R)$ is due to the SSL assumption (Equation 8). In particular,
$\text{pen}(R,\theta)=\sum_{j\in[p]}\left(\lambda_{1}||R(\cdot,j)||_{1}-\sum_{i\in[p]}\log\left(\frac{p_{j}^{*}(0,\theta_{i,j})}{p_{j}^{*}(R(i,j),\theta_{i,j})}\right)\right)$ (13)
where
$p_{j}^{*}(x,\theta_{i,j})=\frac{\theta_{i,j}\psi_{1,j}(x)}{\theta_{i,j}\psi_{1,j}(x)+(1-\theta_{i,j})\psi_{0,j}(x)}$ (14)
However, practically, as suggested by literature ([8] and [1] for instance), the term $\hat{\theta}_{i,j}^{BE}$ may be approximated by
@@
for all $i$. We therefore interact with $\hat{\theta}_{j}^{ABE}$ instead.
@@ 2.2 Asymptotic Analysis
Write $\beta=R^{\intercal}$. The idea of asymptotic analysis is to provide probabilistic inference towards the limiting behaviour of estimation matrix $\hat{R}^{BE}$ (equivalently thereafter $\hat{\beta}$) and the rank $\hat{r}$. In what follows, we introduce some approximate notions of $\hat{\beta}$ and the associated ranks to enable the adaption towards some existing theories and our proposed asymptotic statements.
@@
Below, we briefly outline the posterior concentration from [8], which leads to the proof of Theorem6.
@@
There exists $M$ independent of $p$, $K$ not increasing with $p$ such that as $p,T\to\infty$,
$\forall i,\ \ \sup_{\beta_{i}}\mathbb{E}_{\beta_{i}}\left[\mathbb{P}_{\cdot|D,\theta}\left(\hat{\beta}_{i}:||\beta_{i}-\hat{\beta}_{i}||_{1}>qM\sqrt{\frac{\log p}{T}}\right)\right]$ $\to 0$ (17)
$\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(\hat{\beta}:\gamma(\hat{\beta})>q(1+K))]$ $\to 0$ (18)
The first limit is exactly the same as what has been proved in [8], whereas the second limit can be proved with the same technique, but at a matrix level.
@@
Let $q$ be independent of $p$ and let $p,T$ satisfy $\frac{\log p}{T}>\delta^{2}(qM)^{-2}$ eventually as $p,T\to\infty$
@@
Let $B_{j}:=\{\hat{\beta}:\gamma(\hat{\beta})=j\}$ and $C:=\{\hat{\beta}:\gamma(\hat{\beta})>q(1+K)\}$. Observe
$\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(A)]\leq\sum_{j=0}^{q(1+K)}\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(B_{j}\cap A)]+\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(C\cap A)]$
Now, it is clear from Equation 18 of lemma5 that $\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(C\cap A)]\to 0$ so it remains to prove $\sup_{\beta}\mathbb{P}_{\cdot|D,\theta}(B_{j}\cap A)\to 0$ for all $j\leq q(1+K)$.
@@
    When $j\leq q(1+K)$ and $j\neq q$, consider $\hat{\beta}\in B_{j}$. Observe that there exists $i$ such that $\gamma_{i}(\hat{\beta})\neq||\beta_{i}||_{0}$. Henceforth by construction, $||\tilde{\beta}_{i}-\beta_{i}||_{1}>\delta$. Now, let set $I=\{i:\gamma_{i}(\hat{\beta})\neq||\beta_{i}||_{0}\}$, where cardinality $|I|\leq q(2+K)$. Observe
    $\mathbb{P}_{\cdot|D,\theta}(B_{j}\cap A)\leq\mathbb{P}_{\cdot|D,\theta}(B_{j})\leq q(2+K)\max_{i\in[p]}\mathbb{P}_{\cdot|D,\theta}(\{\hat{\beta}_{i}:||\tilde{\beta}_{i}-\beta_{i}||_{1}>\delta\})$ (19)
    By lemma5, we know $\forall i,\mathbb{P}_{\cdot|D,\theta}(\{\hat{\beta}_{i}:||\tilde{\beta}_{i}-\beta_{i}||_{1}>\delta\})\to 0$ which completes the proof.
@@
    $\mathbb{P}_{\cdot|D,\theta}(B_{q}\cap A)\leq q\max_{i\in I}\mathbb{P}_{\cdot|D,\theta}(||\tilde{\beta}_{i}-\beta_{i}||_{1}>\delta)\to 0$ (20)

### 12 Mar 2024 v2

@@ 1 Introduction
Let $Y$ denote a $p$-dimensional multivariate time series, with $Y_{t}\in\mathbb{R}^{p}$ being observed at time $t\in[T]$, and $Y_{t,j}\in\mathbb{R}$ denoting observation of $j$-th variable at time $t$. We write $Y(\cdot,j)\in\mathbb{R}^{1\times T}$ to denote the $j$-th variable of vector $Y$, which is a univariate time series. Stationarity of $Y(\cdot,j)$ is considered as the second-order stationarity. Precisely, we say $Y(\cdot,j)$ is stationary if the first moment being finite and independent of time and that the auto-covariance function being finite and independent of time. We denote $Y(\cdot,j)\sim I(0)$ in this scenario. When $Y(\cdot,j)\sim I(0)$ for all $j\in[p]$, we write $Y\sim I(0)$. When $Y(\cdot,j)$ is non-stationary but when $\Delta Y(\cdot,j)$ defined by $\Delta Y(t,j):=Y(t,j)-Y(t-1,j)$ is stationary, we write $Y(\cdot,j)\sim I(1)$. Particularly, when $Y(\cdot,j)\sim I(1)$ for all $j\in[p]$, we write $Y\sim I(1)$.
@@ 1.1 VECM and a motivation for High Dimensional VECM
The idea of high dimensional VECM falls in the trendy literature surrounding ‘big data’, in which we have a large amount of variables of interest ($p$) and a substantial amount of observations taken over time ($T$). In a traditional statistical setting, we require more observations than unknown parameters — in the setting of Equation 2 we need $T>p^{2}+\frac{p(p+1)}{2}$. However, that often means we need a greater amount of observations than the data actually have, in order to estimate the equation. This is particularly concerning in light of change-points or other potential changes of the data over time, which naturally constrains the size of $T$. High dimensional VECM is therefore motivated in the regime where $p,T$ are large but $T\leq p^{2}$. Certainly, for estimating the non-sparse coordinates of the matrix $\Pi$, it is reasonable to impose a feasibility assumption where $T>pr$.
@@ A QR decomposition is then proposed, taking the form of
$\tilde{\Pi}=\tilde{R}^{\intercal}\tilde{S}^{\intercal}$ (4)
where $\tilde{R}$ is a p-dimensional upper triangular matrix and $\tilde{S}\in\mathbb{R}^{p\times p}$ is orthogonal. [7] proceed with group-lasso regression to find a low-rank $\hat{R}$ based on $\tilde{R}$, however, we opt for a different route to investigate the true rank of $\Pi$. An important takeaway from this method is that the determination of rank $r$ can be written as $\tilde{r}=rank(\tilde{R})=rank(\tilde{\Pi})$, and subsequently $\hat{r}=rank(\hat{R})=rank(\hat{R}\tilde{S}^{\intercal})$ for an alternative estimator $\hat{R}$ of the true $R$. The rank collected is purely the number of non-zero columns of $R$, as all columns of $R$ are linearly independent, due to the nature of the QR decomposition (such as via Gram–Schmidt process).
The SSL literature mainly concerns scalar response multivariate-regression, i.e. the type taking the form of $y_{i}=x_{i}^{\intercal}\beta+\varepsilon_{i}$ for each index $i$, with $x_{i}$ and $\beta$ being $p-$dimensional vectors and $y_{i}$ being an one-dimensional observation, while $\varepsilon_{i}$ is an iid draw from a Gaussian distribution with zero mean and unknown variance. $\beta_{0}$ is denoted as the true value of the $\beta$ parameter. SSL can be thought of as a mixing distribution being imposed on the parameter $\beta$, conditional on certain hierarchical Bayes parameters $\gamma$, which is then conditionally Bernoulli based on $\theta$, which denotes prior expected fraction of non-zero entries of the true $\beta_{0}$. A posterior mode estimator then aims to maximise the posterior distribution $\pi(\beta|\theta)$. Write $\hat{\beta}=\operatorname*{argmax}\pi(\beta|\theta)$. In [9], $\theta\sim\cal B$ (a Beta distribution on $\theta$) was proposed to enable further inference on the distribution of the estimator conditional on $\theta$, denoted $\pi(\hat{\beta}|\theta)$, and thereafter the conditional expectation $\mathbb{E}[\theta|\hat{\beta}]$. In [10], further analyses were drawn on the asymptotic behaviour of $\hat{\beta}$ such as bounding the limiting posterior probability based on the true $\beta=\beta_{0}$ with data $D$ and $\theta$, namely $\sup_{\beta_{0}}\mathbb{E}_{\beta_{0}}\left[\mathbb{P}_{\cdot|D,\theta}[\hat{\beta}:||\hat{\beta}-\beta_{0}||_{1}>\varepsilon]\right]$ with $\varepsilon$ depending on the number of observation $n$ and features dimension $p$. [3] is another key source of literature that provide asymptotic inference on this type of Bayesian system.
@@ 2 Methodology
2.1 Simple SSL distribution on decomposed VECM
We first introduce the SSL distribution, in both univariate and multivariate contexts, then apply it to the VECM decomposition.
@@ 2.1 Simple SSL distribution on decomposed VECM
Let $X\in\mathbb{R}$ be a random variable. A Spike-and-Slab Lasso (SSL) pior distribution on $X$ is defined as
@@
We write $X|\gamma\sim SSL(\gamma,\lambda_{0},\lambda_{1})$. Let $Y\in\mathbb{R}^{p}$ be a random variable. We write $Y|\gamma\sim SSL(\gamma,\lambda_{0},\lambda_{1})$ if for all entries $j\in[p]$, $Y_{j}|\gamma\sim SSL(\gamma,\lambda_{0},\lambda_{1})$iid
We may interpret $X|\gamma$ as a mixture of two Laplace distribution, with $\psi_{0}$ being spiky at around 0 and $\psi_{1}$ less spiky. $Y|\gamma$ being henceforth a vectorised version of such a distribution, with identical probabilities being attributed.
Recall the Equation 3 and Equation 4 composition of the pre-estimate of $\Pi$. In what follows, we attribute an SSL distribution on the decomposed cointegration matrix in VECM. We use the notation $A_{t}:=\Delta Y_{t}$ and $B_{t}:=\tilde{S}^{\intercal}Y_{t-1}$, with centralisation taken place from $A_{t}$ to $\tilde{A}_{t}$ and normalisation from $B_{t}$ to $\tilde{B}_{t}$ (see appendix A.1 for granular details of data processing).
Definition 2 (Simple SSL distribution on decomposed VECM)
@@
Let $\tilde{A}_{t},\tilde{B}_{t}\in\mathbb{R}^{p}$ for all $t\in[T]$. Let $\theta\in(0,1)^{p}$ be a fixed hyperparameter. A simple SSL distribution on decomposed VECM takes the form of
$\tilde{A}_{t}|\tilde{B}_{t},R,\sigma^{2}$ $\sim N(R^{\intercal}\tilde{B}_{t},\sigma^{2}I_{p}),\ \ \pi(\sigma^{2})\propto\sigma^{-2}$ $\forall t\in[T]$ (7)
$R(\cdot,j)|\gamma_{j}$ $\sim SSL(\gamma_{j},\lambda_{0,j},\lambda_{1}),\ \ \gamma_{j}|\theta_{j}\sim Bernoulli(\theta_{j})$ $\forall j\in[p]$ (8)
In the above model, we have data $\{A_{t},B_{t}\}_{t\in[T]}$ and modelling parameters $R,\sigma$ while $\gamma$ can be marginalised due to their conjugate distributions. As a notational remark: the $\psi_{0,j}(\cdot)$ and $\psi_{1}(\cdot)$ functions are defined accordingly to the $\lambda_{0,j},\lambda_{1}$ as part of the SSL distribution.
@@
Definition 3 (Bayes Estimator)
@@
Let $D$ be a collection of time series data with initialisation $D_{0}$ and $\phi$ be modelling parameters living in the space $\Phi$. If the posterior $\pi(\phi|D,D_{0})$ is well defined over $\phi\in\Phi$ and it attains its unique maximiser, then the Bayes Estimator is the maximiser
$\hat{\phi}^{BE}:=\operatorname*{argmax}_{\phi\in\Phi}\pi(\phi|D,D_{0})$
Under the above setting, we note that $\phi=(R,\sigma)$ and $\Phi=\mathbb{R}^{p\times p}\times\mathbb{R}_{>0}$. We can derive the Bayes Estimator into the followings: define $e_{t}:\mathbb{R}^{p\times p}\to\mathbb{R}^{p}$ by $e_{t}(R)=\tilde{A}_{t}-R^{\intercal}\tilde{B}_{t}$ for all $t$,
$(\hat{R}^{BE},\hat{\sigma}^{BE})=$ $\operatorname*{argmin}_{R\in\mathbb{R}^{p\times p},\sigma\in\mathbb{R}_{>0}}\sum_{t\in[T]}\frac{||{e_{t}}(R)||_{2}^{2}}{2\sigma^{2}}+(pT+2)\log(\sigma)+\text{pen}(R,\theta)$ (9)
$\text{pen}(R,\theta):=$ $\sum_{j\in[p]}\left(\lambda_{1}||R(\cdot,j)||_{1}+\sum_{i\in[p]}\log\left(p_{j}^{*}(R(i,j),\theta_{j})\right)\right)$ (10)
$p_{j}^{*}(x,\theta_{j}):=$ $\frac{\theta_{j}\psi_{1}(x)}{\theta_{j}\psi_{1}(x)+(1-\theta_{j})\psi_{0,j}(x)}$ (11)
Derivation details are at appendix A.2. In essence, the first two terms of Equation 9 are due to Gaussian assumption (Equation 7) with Jeffery prior on the variance. The penalty term $\text{pen}(R,\theta)$ is due to the SSL assumption (Equation 8).
The consequential estimator takes the form of a penalised least square where a $||\cdot||_{1}$ norm is applied to the entire matrix $R$, together with some extra terms measured by $p_{j}^{*}(\cdot,\cdot)$.
2.2 Hierarchical Bayes Modelling of SSL on VECM
The full model utilises the simple SSL distribution as the core, but addes a further distribution on $\theta$ to enrich the hierarchical architecture. Furthermore, the Gaussian assumption of the data is generalised by replacing Equation 7 with the conditional variance noted as $\sigma_{j}^{2}$ for each entry $\tilde{A}_{t,j}$. Formal definition are set as below.
Definition 4 (Hierarchical Bayes Model of SSL distribution on decomposed VECM)
[…]
Let $\tilde{A}_{t},\tilde{B}_{t}\in\mathbb{R}^{p}$ for all $t\in[T]$. Let $a,b\in\mathbb{R}_{>0}^{p}$ be the parameters of prior distribution of $\theta$. A Hierarchical Bayes Model of SSL distribution on decomposed VECM takes the form of
$\tilde{A}_{t,j}|\tilde{B}_{t},R,\sigma_{j}^{2}$ $\sim N(R(\cdot,j)\tilde{B}_{t},\sigma_{j}^{2}),\ \ \pi(\sigma_{j}^{2})\propto\sigma_{j}^{-2}$ $\forall t\in[T],j\in[p]$ (12)
$R(\cdot,j)|\gamma_{j}$ $\sim SSL(\gamma_{j},\lambda_{0,j},\lambda_{1}),\ \ \gamma_{j}|\theta_{j}\sim Bernoulli(\theta_{j}),\ \ \ \theta_{j}\sim{\cal B}(a_{j},b_{j})$ $\forall j\in[p]$ (13)
The Bayes Estimator of this alternative model is derived in appendix A.3, which takes the following form: define vector ${\epsilon}(j)\in\mathbb{R}^{T}$ by $({\epsilon}(j))_{t}=\frac{(e_{t}(R))_{j}}{\sqrt{2}\sigma_{j}}$ for all entry $j$, then
$(\hat{R}^{BE},(\hat{\sigma}_{j}^{BE})_{j\in[p]})=\operatorname*{argmin}_{R\in\mathbb{R}^{p\times p},({\sigma}_{j})_{j\in[p]}\in\mathbb{R}^{p}_{>0}}\sum_{j\in[p]}\left({||{\epsilon}(j)||_{2}^{2}}+(T+2)\log(\sigma_{j})+\lambda_{1}||R(\cdot,j)||_{1}-\log\left(p^{**}_{j}(R,\theta)\right)\right)$ (14)
with $p^{**}_{j}(\cdot,\cdot)$ being defined as
$p^{**}_{j}(R,\theta):=\int_{(0,1)}\prod_{i\in[p]}\frac{\theta_{j}\pi(\theta_{j})}{p_{j}^{*}(R(i,j),\theta_{j})}d\theta_{j}$ (15)
Further manipulation of the function $p^{**}$ may be extended to interact with, at an approximated level, the posterior $\mathbb{E}[\theta_{j}|\hat{R}^{BE}(\cdot,j)]$. See [10] for further discussions.
@@ 2.3 Asymptotic Analysis
Write $\beta=R^{\intercal}$. The idea of asymptotic analysis is to provide probabilistic inference towards the limiting behaviour of the estimation matrix $\hat{R}^{BE}$ (equivalently thereafter $\hat{\beta}$) in Equation 14 and the rank $\hat{r}$, which is defined as the number of non-zero columns of $\hat{R}^{BE}$. In what follows, we introduce some approximate notions of $\hat{\beta}$ and the associated ranks to enable the adaption towards some existing theories and our proposed asymptotic statements.
@@
Below, we briefly outline the posterior concentration, mainly from [10], which leads to the proof of Theorem8. The notation $\mathbb{P}_{\cdot|D,\theta}(\cdot)$ denotes the posterior concentration given data $D$ and true parameter $\theta$.
@@
Write $q_{i}:=||\beta_{i}||_{0}$ and let $q^{M}:=\max_{i\in[p]}q_{i}$. There exists constants $M_{i}>0$ for all $i\in[p]$, and $K>0$ decreasing with $p$ such that as $p,T\to\infty$,
$\forall i,\ \ \sup_{\beta_{i}}\mathbb{E}_{\beta_{i}}\left[\mathbb{P}_{\cdot|D,\theta}\left(\hat{\beta}_{i}:||\beta_{i}-\hat{\beta}_{i}||_{1}>q_{i}M_{i}\sqrt{\frac{\log p}{T}}\right)\right]$ $\to 0$ (17)
$\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(\hat{\beta}:\gamma(\hat{\beta})>rq^{M}(1+K))]$ $\to 0$ (18)
Proof.
The first limit (Equation 17) is exactly the same as what has been proved in [10, Theorem 8], as in our setting, $\beta_{j}|\gamma_{j}$ is SSL-distributed.
The second limit (Equation 18) can be proved with the help of Theorem 7 in [10], which gives a row-level sparsity criteria as follows: there exists $K_{i}>0$ decreasing with $p$ such that
$\sup_{\beta_{i}}\mathbb{E}_{\beta_{i}}[\mathbb{P}_{\cdot|D,\theta}(\hat{\beta}_{i}:\gamma_{i}(\hat{\beta})>q(1+K_{i}))]\to 0\ \ \text{ as }p,T\to\infty$ (19)
Consider the set $I:=\{i\in[p]:\beta_{i}=0\}$. Then by Equation 17, $\sup_{\beta_{i}}\mathbb{E}_{\beta_{i}}\left[\mathbb{P}_{\cdot|D,\theta}\left(\hat{\beta}_{i}:\hat{\beta}_{i}\neq 0\right)\right]\to 0$. Therefore
$\sup_{\beta}\mathbb{E}_{\beta}\left[\mathbb{P}_{\cdot|D,\theta}\left(\hat{\beta}:\gamma_{i}(\hat{\beta})>0\right)\right]\to 0\ \ \ \forall i\in I$ (20)
For $i\notin I$, we pick $K=\max_{i\in[p]\setminus I}K_{i}$. Because of Equation 19, we have
$\sup_{\beta}\mathbb{E}_{\beta}\left[\mathbb{P}_{\cdot|D,\theta}\left(\hat{\beta}:\sum_{i\in[p]\setminus I}\gamma_{i}(\hat{\beta})>\sum_{i\in[p]\setminus I}q(1+K_{i})\right)\right]\to 0$ (21)
By the construction of $r$, we have $|[p]\setminus I|=r$ therefore $rq^{M}(1+K)\geq\sum_{i\in[p]\setminus I}q(1+K_{i})$. This combines with Equation 20 and Equation 21 leads to Equation 18. ∎
@@
Let $r,q^{M}$ be independent of $p$ and let $\delta_{i},p,T$ satisfy $\frac{\log p}{T}\geq\delta_{i}^{2}(qM)^{-2}$ eventually as $p,T\to\infty$, for all $i\in[p]$. Then
@@
Write $q^{\prime}:=rq^{M}(1+K)$. Let $A:=\{\hat{\beta}:\tilde{r}(\hat{\beta})\neq r\}$, $B_{j}:=\{\hat{\beta}:\gamma(\hat{\beta})=j\}\forall j\leq q^{\prime}$ and $C:=\{\hat{\beta}:\gamma(\hat{\beta})>q^{\prime}\}$. Observe
$\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(A)]\leq\sum_{j=0}^{q^{\prime}}\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(B_{j}\cap A)]+\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(C\cap A)]$
Now, it is clear from Equation 18 of lemma7 that $\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(C\cap A)]\leq\sup_{\beta}\mathbb{E}_{\beta}[\mathbb{P}_{\cdot|D,\theta}(C)]\to 0$ so it remains to prove $\sup_{\beta}\mathbb{P}_{\cdot|D,\theta}(B_{j}\cap A)\to 0$ for all $j\leq q^{\prime}$.
@@
    When $j\leq q^{\prime}$ and $j\neq q$, consider $\hat{\beta}\in B_{j}$. Observe that there exists $i$ such that $\gamma_{i}(\hat{\beta})\neq||\beta_{i}||_{0}$. Henceforth by construction, $||\tilde{\beta}_{i}-\beta_{i}||_{1}>\delta$. Now, let set $I=\{i:\gamma_{i}(\hat{\beta})\neq||\beta_{i}||_{0}\}$, where cardinality $|I|\leq q+q^{\prime}$ as $j\leq q^{\prime}$. Observe
    $\mathbb{P}_{\cdot|D,\theta}(B_{j}\cap A)\leq\mathbb{P}_{\cdot|D,\theta}(B_{j})\leq(q+q^{\prime})\max_{i\in I}\mathbb{P}_{\cdot|D,\theta}(\{\hat{\beta}_{i}:||\tilde{\beta}_{i}-\beta_{i}||_{1}>\delta_{i}\})$ (22)
    Now notice, for all $i\in I$, by construction, $\tilde{\beta_{i}}=\hat{\beta_{i}}$, and so by lemma7, we know $\mathbb{P}_{\cdot|D,\theta}(\{\hat{\beta}_{i}:||\tilde{\beta}_{i}-\beta_{i}||_{1}>\delta_{i}\})\to 0$ which completes the proof.
@@
    $\mathbb{P}_{\cdot|D,\theta}(B_{q}\cap A)\leq(q+q^{\prime})\max_{i\in I}\mathbb{P}_{\cdot|D,\theta}(\{\hat{\beta}_{i}:||\tilde{\beta}_{i}-\beta_{i}||_{1}>\delta_{i}\})\to 0$ (23)
Theorem8 offers some level of assurance in terms of consistency in rank estimation. In essence, the statement gives a strong probability convergence, globally over $\beta\in\mathbb{R}^{p\times p}$, of the generalised rank $\tilde{r}$ to $r$.